OpenStudy (anonymous):

Can someone explain how to add fractions with different common denominators? e.g. 16/48 + 1/8?

OpenStudy (heisenberg):

So to do operations on fractions, their denominator (the number on the bottom) has to be the same. We can use one basic property of fractions to help this: \[\frac{x}{x} = 1\]

OpenStudy (heisenberg):

We can multiply our fraction by something over itself to get it in a new form: \[\frac{1}{8} * \frac{6}{6} = \frac{6}{48}\]

OpenStudy (anonymous):

right, so you must find a common multiple for the values of the denominators

OpenStudy (heisenberg):

now since they have a common denominator, we can do the operation: \[\frac{16}{8} + \frac{6}{48} = \frac{22}{48}\]

OpenStudy (heisenberg):

OOPS! that should be \[\frac{16}{48}\]

OpenStudy (anonymous):

guys this is really helpful

OpenStudy (heisenberg):

great! ask any more questions you have on the left! if i helped you, please fan me :)

OpenStudy (anonymous):

Thanks guys, I understand better now!

OpenStudy (anonymous):

you didn't simplify 22/48 it should be 11/24

OpenStudy (anonymous):

ya wanna see another way? I use this to get all the fractions out of an equation

OpenStudy (anonymous):

16/48 + 1/8 = x Multiply everything by the first denominator 16 + 48/8 = 48x then the second denominator (repeat this for every fraction) 128 + 48 = 384x then solve 176/384 = x and simplify 11/24 = x